Invited plenary speakers

Albrecht Boettcher (Chemnitz, Germany) Froilán Dopico (Madrid, Spain) Heike Fassbender (Braunschweig, Germany) Luca Gemignani (Pisa, Italy) Leslie Hogben (Ames, Iowa, USA) Erxiong Jiang (Shanghai, China) Daniel Kressner (Zurich, Switzerland) James Nagy (the SIAM Lecture) (Atlanta, Georgia, USA) Juan Manuel Peña (Zaragoza, Spain) Peter Rosenthal (Toronto, Canada) Naomi Shaked-Monderer (Emek Yezreel, Israel) Ilya Spitkovsky (Williamsburg, Virginia, USA) Paul Van Dooren (Louvain, Belgium)

Mini-Symposia

MS1 Combinatorial Matrix Theory

Organizers: Leslie Hogben, Iowa State University, lhogben@iastate.edu Bryan Shader, University of Wyoming, bshader@uwyo.edu

Confirmed participants: Wayne Barrett, Brigham Young University Richard Brualdi, University of Wisconsin Luz DeAlba, Drake University Danny Hershkowitz, Technion Hein van der Holst, Eindhoven University of Technology Steve Kirkland, University of Regina Dale Olesky, University of Victoria Olga Pryporova, Iowa State University Michael Tsatsomeros, Washington State University Kevin Vander Meulen, Redeemer College

MS2 Eigenproblems: Theory and computation

Organizer: Julio Moro, Universidad Carlos III de Madrid, jmoro@math.uc3m.es

Confirmed participants: Tom Bella Department of Mathematics University of Connecticut e-mail:

bella@math.uconn.edu

Steve Cox Dept. of Computational and Applied Mathematics Rice University e-mail: cox@rice.edu

Michael Karow Institut fr Mathematik Technische Universitt Berlin email: karow@math.tu-berlin.de

Niloufer Mackey Department of Mathematics Western Michigan University e-mail: nil.mackey@wmich.edu

Karl Meerbergen Katholieke Universiteit Leuven Department of Computer Science e-mail: Karl.Meerbergen@cs.kuleuven.be

Julio Moro Departamento de Matemticas Universidad Carlos III de Madrid e-mail: jmoro@math.uc3m.es

Vadim Olshevsky Department of Mathematics University of Connecticut e-mail: olshevsky@math.uconn.edu

Bor Plestenjak Department of Mathematics, University of Ljubljana email: Bor.Plestenjak@fmf.uni-lj.si

MS3 Implementation and Application issues in regularizing least squares and total least squares.

Organizer: Rosie Renaut, Arizona State University, renaut@asu.edu

Confirmed participants: John Bardlsey bardsleyj@mso.umt.edu Truncation Rules for Iterative Deblurring Methods Antonio Arico, arico@unica.it Signal and Image regularization via antireflective transform Jodi Mead, Boise State University, jmead@boisestate.edu, and Rosemary

Renaut

Calculating Weights in Least Squares Estimation Using the Chi-squared Method

Rosemary A Renaut, Arizona State University, and Jodi Mead

A Newton Iteration for estimating the regularizing parameter for least squares

Iveta Hnětynková, Faculty of Mathematics and Physics, Charles University in Prague, and Institute of Computer Science, Academy of Sciences, Czech Republic. hnetynkova@cs.cas.cz

On solvability of total least squares problem Wolfgang Stefan, wolfgang.stefan@gmail.com, and Anne Gelb, Regularizing Least Squares with the Concentration Method Julianne Chung, James Nagy, Emory University, Lanczos Hybrid Regularization for Image Processing Applications Bert W. Rust, bwr@nist.gov, and Dianne P. O’Leary, A Truncated Singular Component Method for Ill-Posed Problems

MS4 Linear Algebra Education

Organizers: Maria Trigueros, Instituto Tecnologico Autonomo de Mexico, trigue@itam.mx Asuman Oktac, Centro de Investigacion y Estudios Avanzados, oktac@mercurio.red.cinvestav.mx

Confirmed participants:

1. Hamide Dogan-Dunlap (USA/Turkey): Thinking Modes Revealed on Students’ Responses from an Assignment on Linear Independence.

2. Jacqueline Klasa-Bompoint (Canada): Use of non-standard bases in the study of conics, Illustrations with Cabri

3. Denis Tanguay (Canada)

4. Laurel Cooley (group of presenters) (USA): Integrating learning theories and application-based modules in teaching linear algebra

5. Mirko Maracci (Italy)

6. Michelle Zandieh (group) (USA)

7. Ana Moura Santos (Portugal)

8. Gustavo Preciado (group) Mexico: Use of models in the teaching of linear algebra

9. Asuman Okta and Maria Trigueros (group) (Mexico): On the learning of the concept of base

MS5 Linear Algebra in Model Reduction

Organizer: Peter Benner, Chemnitz University of Technology

Confirmed participants:

1. Serkan Gugercin (Virginia Tech, Blacksburg, USA; gugercin@math.vt.edu): A Krylov-Based Descent Approach for the Optimal H2 Model Reduction

of Large-Scale Dynamical Systems

2. Paul Van Dooren (Universite Catholique Louvain-La-Neuve, Belgium; paul.vandooren@uclouvain.be):

H2 Approximation and Tangential Rational Interpolation

3. Christopher Beattie (Virginia Tech, Blacksburg, USA; beattie@vt.edu): Interpolatory Projection Methods for Parameterized Model Reduction

4. Lihong Feng (Chemnitz University of Technology, Germany/Fudan University Shang- hai, China; lhfeng@fudan.edu.cn):

Parametric Model Reduction for Systems with Coupled Parameters

5. Roland Freund (University of California, Davis, USA; freund@math.ucdavis.edu): The Effects of Deflation in Projection-Based Order Reduction

6. Ulrike Baur (Chemnitz University of Technology, Germany; baur@mathematik.tuchemnitz.de):

Model Reduction for unstable Systems based on Hierarchical Matrix Arithmetic

7. Tobias Damm (TU Kaiserslautern, Germany; damm@mathematik.unikl.de):

Algebraic Gramians and Model Reduction for Dient System Classes

8. Peter Benner (Chemnitz University of Technology, Germany; benner@mathematik.tu-chemnitz.de): Balancing-Related Model Reduction for Large-Scale Unstable Systems

MS6 Matrix functions and matrix equations

Organizers: Peter Lancaster, University of Calgary, lancaste@ucalgary.ca Chun-Hua Guo, University of Regina, chguo@math.uregina.ca

Confirmed participants: Amir Amiraslani, University of Calgary, amiram@math.ucalgary.ca Peter Benner, TU Chemnitz, benner@mathematik.tu-chemnitz.de Dario Bini, University of Pisa, bini@dm.unipi.it Lyonell Boulton, HeriotWatt University, Edinburgh, lyonell@ma.hw.ac.uk Heike Fassbender, TU Braunschweig, h.fassbender@tu-bs.de Chun-Hua Guo, University of Regina, chguo@math.uregina.ca Bruno Iannazzo, University of Pisa, iannazzo@mail.dm.unipi.it Peter Lancaster, University of Calgary, lancaste@ucalgary.ca Beatrice Meini, University of Pisa, meini@dm.unipi.it Shu-Fang Xu, Peking University, xsf@pku.edu.cn

MS7 Max Algebra

Organizers: Hans Schneider, University of Wisconsin Peter Butkovic, University of Birmingham, UK

Confirmed participants: Stephane Gaubert, INRIA Paris, France stephane.gaubert@inria.fr Using max-plus eigenvalues to bound the roots of a polynomial (with

Marianne Akian (INRIA) Adrien Brandejsky (ENS Cachan))

Jean Jacques Loiseau,University Nantes, France jean-jacques.loiseau@irccyn.ec-nantes.fr

Robust stability of positive difference equations (with Michael Di Loreto) Sergei Sergeev, University of Birmingham, UK sergiej@gmail.com Cellular-type conditions for two-matrix eigenproblem Ivan Singer, Institute of Mathematics, Bucharest, Romania ivan.singer@imar.ro Max-min convexity Bill McEneaney, University of California, San Diego, USA wmceneaney@ucsd.edu An Idempotent Semiring Underlying New Fundamental Solutions of Dif-

ferential Equations with Quadratic Nonlinearities Martin Gavalec, University Hradec Kralove, Czech Rep martin.gavalec@uhk.cz Permuted max-min eigenvector problem (with J.Plavka) Glenn Merlet, University Paris-Dauphine, France glenn.merlet@gmail.com Ergodic theorems for products of random matrices in the max-plus alge-

bra Jean-Louis Boimond, University Angers, France boimond@istia.univ-angers.fr On Just In Time Controller Design in Max-Plus Algebra Hans Schneider, University of Wisconsin, Madison, USA hans@math.wisc.edu TBA Peter Butkovic, University of Birmingham, UK p.butkovic@bham.ac.uk On the permuted max-algebraic eigenvector problem Jan Plavka, Technical University Kosice, Slovakia jan.plavka@tuke.sk On the robustness of matrices in max-min algebra Peter Szabo, Technical University Kosice, Slovakia peter.szabo@tuke.sk A Note on Partitions and Triangular Toeplitz matrices Monika Molnarova, Technical University Kosice, Slovakia monika.molnarova@tuke.sk Orbit periods with respect to Monge matrices in max-plus algebra

MS8 Nonnegative and eventually nonnegative matrices

Organizer: Judi McDonald, Washington State University, jmcdonald@math.wsu.edu

Confirmed participants: Minerva Catral, University of Victoria, mcatral@math.uvic.ca Sherod Eubanks, Washington State University, eubanks@math.wsu.edu Daniel Hershkowitz, Techion, hershkow@techunix.technion.ac.il Thomas Laffey, University College Dublin, Thomas.Laffey@ucd.ie Judi McDonald, Washington State University, jmcdonald@math.wsu.edu Michael Neumann, University of Conneticut, neumann@math.uconn.edu DeAnne Morris, Washington State University, dmorris@math.wsu.edu Helena Smigoc, Univeristy College Dublin, helena.smigoc@fmf.uni-lj.si Michael Tsatsomeros, Washington State University, tsat@math.wsu.edu